OpenStudy (anonymous):

let the width of rectangle FRAN equal w. The rectangle is 4cm less than the width. In the rectangle IREN the width is 7 more than w and the length is 2 less than 3 times w. if their perimeters are equal, find w.

OpenStudy (turingtest):

Your question doesn't make sense: "The rectangle is 4cm less than the width."?

OpenStudy (anonymous):

sorry typo it was the LENGTH of the rectangle is 4cm less than the width

OpenStudy (turingtest):

perimeter of fran we call A= 2(w+L)=2[w+(w+4)]=4w+8 perimeter of iren we call B= 2(w+L)=2[(w+7)+(3w-2)]=8w-10 A=B 4w+8=8w-10 4w=18 w=4.5 pretty sure I got it